OpenStudy (anonymous):

A quiz consists of 4 true-false questions. What is the probability of getting 100% on the quiz by randomly guessing the answer to all 4 questions? 1/2 2/5 1/16

OpenStudy (anonymous):

p true = 1/2 with 4 questions right: => P ( 4 true ) = ( 1/2) ^4 =